OpenStudy (anonymous):

3sqroot6 + 5sqroot2 over 4sqroot6 - 3sqroot2

OpenStudy (anonymous):

well do a graph equation and do a qubic square root and do the order of operations and a inverse operation

OpenStudy (anonymous):

Is this the equation\[3\sqrt{6}+\frac{5\sqrt{2} }{4\sqrt{6} }-3\sqrt{2}\]?

OpenStudy (anonymous):

yep:)

OpenStudy (anonymous):

so what are you required to do with it?

OpenStudy (anonymous):

Not much simplification can be done